# Generalised height pairings and the Albanese kernel

### v1

@@ 2.8. Generalised heights and quadratic Chabauty
$X(K)\to M(\operatorname{Gal}(K);\mathbb{Q}_{p},V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$
@@
hence we obtain a map from $H^{1}(K,U_{2})$ to $H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2}/\mathbb{Q}_{p}(1)))$ as desired. In the language of mixed extensions, this sends a $U_{2}$-torsor $P$ to the twist of $E_{2}(X;b,z)$ by $P$.
@@ 3. The Albanese kernel and mixed extensions from algebraic cycles
3.1. The relation with mixed extensions
@@
[…]
@@ 3.2. A space of motivic mixed extensions
Define $\mathcal{M}(X\times X)$ to be the set of tuples $(Z_{1},Z_{2},(C_{i},D_{i}))$, where
@@
    $D_{i}$ is a principal divisor on $D_{i}$.
@@
modulo the equivalence relation that $(Z_{1},Z_{2},(C_{i},D_{i}))$ is equivalent to $(Z^{\prime}_{1},Z^{\prime}_{2},(C^{\prime}_{i},D^{\prime}_{i})$ if $Z_{i}\sim Z^{\prime}_{i}$ in $\mathrm{CH}^{1}(X)_{0}$, and
$\sum(Z_{1}\times X,Z_{2}-Z^{\prime}_{2})+\sum(Z^{\prime}_{1},Z^{\prime}_{2}-Z_{2})+\sum(C_{i},D_{i})-\sum(C^{\prime}_{i},D^{\prime}_{i})$
lies in the image of $K^{M}_{2}(X^{2})$ under the natural map
$K^{M}_{2}(X^{2})\to\oplus_{C\in X^{(1)}}\mathbb{Q}(C)^{\times}\to\oplus_{C\in X^{(1)}}\mathrm{Div}(C),$
where $X^{(p)}$ denote the set of codimension $p$ points of the Zariski topological space $X$. We obtain a map
@@
which lands in the kernel of the map to $T(X\times X)$. Let $H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2}))_{s}\subset H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2})$ denote the set of mixed extensions whose image in $H^{1}(K,V^{*}\otimes V^{\otimes 2})$ lies in the image of $H^{1}(K,V)$ via the left multiplication map
$V\to V^{*}\otimes V^{\otimes 2}.$
Let $|C|:=\cup C_{i}$. Let $N\subset V^{\otimes 2}$ be the the first step of the coniveau filtration, i.e. the subspace generated by $H^{2}_{C}(X^{2},\mathbb{Q}_{p}(2))$ for curves $C\subset X^{2}$.
@@
[…]
@@
$0\to H^{1}_{\mathrm{et}}(X^{2}-|Z_{1}|)\otimes H^{1}(X)/N\to E\to\mathbb{Q}_{p}\to 0$
@@ 3.3. A generalisation
where $\pi_{i}$ is the projection $X_{1}\times X_{2}\to X_{i}$. Then Beilinson’s conjecture implies that this pairing, restricted to $\mathrm{CH}^{p_{1}}(X_{1})_{0}\times\mathrm{CH}^{p_{2}}(X_{2})_{0}$, is torsion. If $\pi_{1}^{*}Z_{1}\cdot\pi_{2}^{*}Z_{2}$ is zero, then we get a codimension $p_{1}+p_{2}-1$ subvariety $W$ and a rational function on $W$ such that $\pi_{1}^{*}Z_{1}\cdot\pi_{2}^{*}Z_{2}=\mathrm{div}(f)$. Define $\mathcal{M}_{p_{1},p_{2}}(X_{1},X_{2})$ to be the set of quadruples $(Z_{1},Z_{2},(W_{i},D_{i}))$ where $Z_{1}$ is in $Z^{p_{1}}(X_{1})$, $Z_{2}$ is in $Z^{p_{2}}(X_{2})$, $W_{i}$ is a codimension $p_{1}+p_{2}-1$ subvariety of $X_{1}\times X_{2}$, $D_{i}$ is a principal divisor on $W_{i}$ such that $\sum_{i}D_{i}=Z_{1}\boxtimes Z_{2}$, modulo the equivalence relation
$(Z_{1},Z_{2},(W_{i},D_{i}))\sim(Z^{\prime}_{1},Z^{\prime}_{2},(W^{\prime}_{i},D^{\prime}_{i}))$
if $Z_{i}-Z^{\prime}_{i}=\mathrm{div}_{W_{i}}(F_{i})$ for $i=1,2$, and
$\sum(Z_{1}\times W_{2},1\times F_{2})+\sum(Z^{\prime}_{1}\times W_{2},1\times F_{2})+\sum(W_{i},D_{i})-\sum(W^{\prime}_{i},D^{\prime}_{i})$
lies in the image of $\oplus_{x\in(X_{1}\times X_{2})^{(p_{1}+p_{2}-2)}}K^{M}_{2}(K(x))$. We get an exact sequence of pointed sets
@@
where $H^{i}:=H^{2d_{i}-1}(X,\mathbb{Q}_{p}(d_{i}))$, and $U(\mathbb{Q}_{p},H_{1},H_{1}\otimes H_{2})_{s}$ is the subgroup of $U(\mathbb{Q}_{p},H_{1},H_{1}\otimes H_{2})$ consisting of automorphisms whose projection to $\mathrm{Hom}(H_{1},H_{1}\otimes H_{2})$ lies in the image of $H_{2}$. These two exact sequences can be related by considering an appropriate subquotient of $H^{2(d_{1}+d_{2}-1)}(X_{1}\times X_{2}-X_{1}\times|Z_{2}|\cup W,\mathbb{Q}_{p}(d_{1}+d_{2}))$, as in Proposition 1.
@@ 5.2. An analogue for the unit equation
$\langle[x]+[y]+[(1-x)/(1-xy)]+[1-xy]+[(1-y)/(1-xy)]\rangle\to\mathbb{Z}[K^{\times}-\{1\}]\stackrel{{\scriptstyle[x]\mapsto x\wedge(1-x)}}{{\longrightarrow}}\wedge^{2}K^{\times}$
@@ 5.4. Circle-valued generalised heights
$\to\mathrm{Ext}^{1}_{\mathbb{R}-\mathrm{MHS}}(\mathbb{R},H^{2d_{1}-1}(X_{1},\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{R}(d_{2})))/\mathrm{reg}\mathrm{CH}^{d_{1}+d_{2}}(X_{1}\times X_{2},1).$
which may be defined as follows. Taking cohomological realisations we obtain a mixed extension of $\mathbb{R}$-mixed Hodge structures with graded pieces $\mathbb{R},H^{2d-1}(X_{1},\mathbb{R}(d_{1}))$ and $H^{2d_{1}-1}(X_{1},\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{R}(d_{2}))$. Mixed extensions of $\mathbb{R}$-mixed Hodge structures were recently studied by Hain [Hai24]. Generalising the usual construction for the height pairing (as in [Sch94]) he shows how one may associated to such a mixed extension an element of $\operatorname{Ext}^{1}_{\mathbb{R}-\mathrm{MHS}}(\mathbb{R},H^{2d_{1}-1}(X_{1},\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{R}(d_{2})))$, called the real period of the mixed extension.
The ‘Tate’ part of $H^{2d_{1}-1}(X_{1},\mathbb{Z}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{Z}(d_{2}))$ should presumably recover the height pairing as defined by Scholl. In general, if the analogy in 5.3 is reasonable, it is natural to ask whether the product constructed is related to the intersection product on arithmetic Chow groups [BG97]
@@
5.5. $p$-descent for Bloch–Kato Selmer groups
In [Dog23], [Dog24] and [Ber25], the problem of bounding the dimension of $H^{1}_{f}(\mathbb{Q},\wedge^{2}V)$ via the construction of a suitable ‘Selmer’ subspace of $H^{1}(\mathcal{Z}_{S},\wedge^{2}J[2])$ is considered. This is somewhat analogous to bounding the rank of an abelian variety via $2$-descent. The analogue of the usual Kummer exact sequence for an abelian variety gives a commutative diagram
@@ 5.6. Other varieties
In the above we used the curves on the surface $X\times X$ to construct mixed extensions. It is natural to wonder whether there is a ‘simpler’ surface (or simpler variety) to work with. In the case of a genus 2 curve $X$, two natural candidates are its Jacobian $J$, and its Kummer variety $K$, since their $H^{2}$ captures the ‘interesting part’ (c.f. Lemma 8) of $H^{2}(X\times X)$. This might give a simpler presentation of the example in section 4.

### v2

@@ 2.8. Generalised heights and quadratic Chabauty
$X(K)\to M_{g}(\operatorname{Gal}(K);\mathbb{Q}_{p},V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$
@@
hence we obtain a map from $H^{1}(K,U_{2})$ to $H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2}/\mathbb{Q}_{p}(1)))$ as desired. In the language of mixed extensions, this sends a $U_{2}$-torsor $P$ to the twist of $E_{2}(X;b)$ by $P$.
@@ 3. The Albanese kernel and mixed extensions from algebraic cycles
The map $\boxtimes$ has the following compatibility with the cup product in Galois cohomology.
@@
$\mathrm{CH}^{1}(X)_{0}\times\mathrm{CH}^{1}(X)_{0}$$T(X^{2})$$H^{1}(K,H^{1}_{\mathrm{et}}(X,\mathbb{Z}_{p}(1)))\times H^{1}(K,H^{1}_{\mathrm{et}}(X,\mathbb{Z}_{p}(1)))$$H^{2}(K,H^{2}_{\mathrm{et}}(X^{2},\mathbb{Z}_{p}(2)))$}
@@ 3.1. A space of motivic mixed extensions
Define $\mathcal{M}(X\times X)$ to be the set of tuples $(Z_{1},Z_{2},\sum(C_{i},D_{i}))$, where
@@
    $D_{i}$ is a principal divisor on $D_{i}$ (so $\sum(C_{i},D_{i})$ is an element of the direct sum, over $C\in(X^{2})^{(1)}$, of $\mathrm{Div}^{0}(C)$).
@@
modulo the equivalence relation generated by the following relations that $(Z_{1},Z_{2},(C_{i},D_{i}))$ is equivalent to $(Z^{\prime}_{1},Z^{\prime}_{2},(C^{\prime}_{i},D^{\prime}_{i})$ if $Z_{i}\sim Z^{\prime}_{i}$ in $\mathrm{CH}^{1}(X)_{0}$, and
$(Z_{1},Z_{2},(C_{i},D_{i}))\sim(Z_{1}+\mathrm{div}(g),Z_{2},\sum(C_{i},D_{i})+(\pi_{2}^{*}Z_{2},\mathrm{div}(g))$
$(Z_{1},Z_{2},(C_{i},D_{i}))\sim(Z_{1},Z_{2}+\mathrm{div}(g),\sum(C_{i},D_{i})+(\pi_{1}^{*}Z_{1},\mathrm{div}(g))$
$(Z_{1},Z_{2},(C_{i},D_{i}))\sim(Z_{1},Z_{2},\sum(C_{i},D_{i})+\partial(\{g_{1},g_{2}\}))$
where $\partial\{g_{1},g_{2}\}$ is the image of $\{g_{1},g_{2}\}\in K^{M}_{2}(K(X^{2}))$ in $\oplus_{C\in X^{(1)}}\mathrm{Div}^{0}(C)$ under the residue map. We obtain a map
@@
which lands in the kernel of the map to $T(X\times X)$. Let $|C|:=\cup C_{i}$. Let $N\subset V^{\otimes 2}$ be the the first step of the coniveau filtration, i.e. the subspace generated by $H^{2}_{C}(X^{2},\mathbb{Q}_{p}(2))$ for curves $C\subset X^{2}$. Let $H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2})/N)_{s}\subset H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2}/N)$ denote the set of mixed extensions whose image in $H^{1}(K,V^{*}\otimes V^{\otimes 2}/N)$ lies in the image of $H^{1}(K,V)$ via the left multiplication map
$V\to V^{*}\otimes V^{\otimes 2}/N$
(note that via this map $V$ is a direct summand of $V^{*}\otimes V^{\otimes 2}/N$).
@@
$\mathrm{CH}^{2}(X^{2},1)$$\mathcal{M}(X^{2})$$\mathrm{CH}^{1}(X)_{0}\times\mathrm{CH}^{1}(X)_{0}$$T(X^{2})$$H^{1}(K,V^{\otimes 2}/N)$$H^{1}(K,U(\mathbb{Q}_{p},V,V^{\otimes 2}))_{s}$$H^{1}(K,V)\times H^{1}(K,V)$$H^{2}(K,V^{\otimes 2}).$}
@@
$0\to H^{1}_{\mathrm{et}}(X^{2}-|Z_{1}|)\otimes H^{1}(X)/N\to E\to\mathbb{Q}_{p}\to 0.$
@@ 3.2. A generalisation
where $\pi_{i}$ is the projection $X_{1}\times X_{2}\to X_{i}$. Then Beilinson’s conjecture implies that this pairing, restricted to $\mathrm{CH}^{p_{1}}(X_{1})_{0}\times\mathrm{CH}^{p_{2}}(X_{2})_{0}$, is torsion. If $\pi_{1}^{*}Z_{1}\cdot\pi_{2}^{*}Z_{2}$ is zero, then we get a codimension $p_{1}+p_{2}-1$ subvariety $W$ and a rational function on $W$ such that $\pi_{1}^{*}Z_{1}\cdot\pi_{2}^{*}Z_{2}=\mathrm{div}(f)$. Define $\mathcal{M}_{p_{1},p_{2}}(X_{1},X_{2})$ to be the set of quadruples $(Z_{1},Z_{2},\sum(W_{i},D_{i}))$ where $Z_{1}$ is in $Z^{p_{1}}(X_{1})$, $Z_{2}$ is in $Z^{p_{2}}(X_{2})$, $W_{i}$ is a codimension $p_{1}+p_{2}-1$ subvariety of $X_{1}\times X_{2}$, $D_{i}$ is a principal divisor on $W_{i}$ such that $\sum_{i}D_{i}=Z_{1}\boxtimes Z_{2}$, modulo the equivalence relation generated by
$(Z_{1},Z_{2},\sum(W_{i},D_{i}))\sim(Z_{1}+\mathrm{div}_{Y_{1}}(g_{1}),Z_{2},\sum(W_{i},D_{i})+(\pi_{1}^{*}Y_{1},\pi_{2}^{*}Z_{2}))$
$(Z_{1},Z_{2},\sum(W_{i},D_{i}))\sim(Z_{1},Z_{2}+\mathrm{div}_{Y_{2}}(g_{2}),\sum(W_{i},D_{i})+(\pi_{2}^{*}Y+2,\pi_{1}^{*}Z_{1}))$
$(Z_{1},Z_{2},\sum(W_{i},D_{i}))\sim(Z_{1},Z_{2},\sum(W_{i},D_{i})+\partial(\{h_{1},h_{2}\}).$
Here $Y_{i}$ is a codimension $p_{i}-1$ subvariety of $X_{i}$, $h_{1}$ and $h_{2}$ are functions on a codimension $p_{1}+p_{2}-2$ subvariety $V$ of $X_{1}\times X_{2}$, and $\partial(\{h_{1},h_{2}\})$ denotes the image of $\{g_{1},g_{2}\}\in K^{M}_{2}(K(V))$ in $\oplus_{x\in X^{p_{1}+p_{2}-1}}\mathrm{Div}(K(x))$ under the residue map. We get an exact sequence of pointed sets
@@
where $H^{i}:=H^{2d_{i}-1}(X_{i},\mathbb{Q}_{p}(d_{i}))$, and $U(\mathbb{Q}_{p},H_{1},H_{1}\otimes H_{2})_{s}$ is the subgroup of $U(\mathbb{Q}_{p},H_{1},H_{1}\otimes H_{2})$ consisting of automorphisms whose projection to $\mathrm{Hom}(H_{1},H_{1}\otimes H_{2})$ lies in the image of $H_{2}$. The relation between the two exact sequences appears to be more complicated than in Proposition 1, because the coniveau filtration on $H^{2(d_{1}+d_{2}-1)}(X_{1}\times X_{2})$ is more complicated than that of $H^{2}$ of a surface.
@@ 5.2. An analogue for the unit equation
$\langle[x]+[y]+[(1-x)/(1-xy)]+[1-xy]+[(1-y)/(1-xy)]|x,y\in K^{\times}-\{1\}\rangle\to\mathbb{Z}[K^{\times}-\{1\}]\stackrel{{\scriptstyle[x]\mapsto x\wedge(1-x)}}{{\longrightarrow}}\wedge^{2}K^{\times}$
@@ 5.4. Circle-valued generalised heights
$\to\mathrm{Ext}^{1}_{\mathbb{R}-\mathrm{MHS}}(\mathbb{R},H^{2d_{1}-1}(X_{1},\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{R}(d_{2})))/(\mathrm{reg}\mathrm{CH}^{d_{1}+d_{2}}(X_{1}\times X_{2},1)+N),$
where $N\subset H^{2d-1}(X,\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X,\mathbb{R}(d_{2}))$ is the image of $H^{2(d_{1}+d_{2}-1)}$ of all codimension $d_{1}+d_{2}-1$ subvarieties, which may be defined as follows. Taking cohomological realisations we obtain a mixed extension of $\mathbb{R}$-mixed Hodge structures with graded pieces $\mathbb{R},H^{2d-1}(X_{1},\mathbb{R}(d_{1}))$ and $H^{2d_{1}-1}(X_{1},\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{R}(d_{2}))/N$. Mixed extensions of $\mathbb{R}$-mixed Hodge structures were recently studied by Hain [Hai24]. Generalising the usual construction for the height pairing (as in [Sch94]) he shows how one may associated to such a mixed extension an element of $\operatorname{Ext}^{1}_{\mathbb{R}-\mathrm{MHS}}(\mathbb{R},H^{2d_{1}-1}(X_{1},\mathbb{R}(d_{1}))\otimes H^{2d_{2}-1}(X_{2},\mathbb{R}(d_{2}))/N)$, called the real period of the mixed extension.
In general, if the analogy in 5.3 is reasonable, it is natural to ask whether the product constructed is related to the intersection product on arithmetic Chow groups [BG97]
@@
$5.$ $p$-descent for Bloch–Kato Selmer groups
In [Dog23], [Dog24] and [Ber25], the problem of bounding the dimension of $H^{1}_{f}(\mathbb{Q},\wedge^{2}V)$ via the construction of a suitable ‘Selmer’ subspace of $H^{1}(\mathbb{Q},\wedge^{2}J[2])$ is considered. This is somewhat analogous to bounding the rank of an abelian variety via $2$-descent. The analogue of the usual Kummer exact sequence for an abelian variety gives a commutative diagram
@@ 5.6. Other varieties
In the above we used the curves on the surface $X\times X$ to construct mixed extensions. It is natural to wonder whether there is a ‘simpler’ surface (or simpler variety) to work with. In the case of a genus 2 curve $X$, two natural candidates are its Jacobian $J$, and its Kummer variety $K$, since their $H^{2}$ captures the ‘interesting part’ (c.f. Lemma 8) of $H^{2}(X\times X)$.
This gives a simpler presentation of the example in section 4 as follows. Instead of working with the genus $8$ curve on $X\times X$, one can instead work with its image in $\mathrm{Jac}(X)$, which is a curve of genus 4. The main computational problem is then to find principal divisors on curves which exhibit the torsion-ness of zero-cycles on $\mathrm{Jac}(X)$. For our genus 4 curve, using the hyperelliptic involution this reduces to finding principal divisors on its image in the Kummer variety which lift to rational points on $\mathrm{Jac}(X)$. Hence recent work of Gazaki and Love [GL23] suggests a systematic way to compute generalised heights, by finding rational curves on the Kummer variety whose pullbacks contain rational points on $\mathrm{Jac}(X)$.